# Deviations from the Nash equilibrium and emergence of tacit collusion in a two-player optimal execution game with reinforcement learning

### v1

@@ 4.2 The moderate noise case
In this setting, the mid-price dynamics is influenced both by the trading of the agents and by the volatility. The first thing that can be noticed in Figure 3 is that the $IS$ per iteration distributes almost only in the second, third and fourth quadrant of the plot. Moreover, the distribution of the points suggests an inverse relation between the $IS$s of the two agents, i.e. a lower $IS$ of an agent is typically associated with a worsening of the $IS$ of the other agent. The centroids distribute accordingly. In fact, the centroids that lie in the collusive area of the graph are as numerous as those that lie outside of that area on either the second or fourth quadrant. It can be seen that, outside the green rectangle, the agents tend to behave in a predatory way, meaning that one agent has consistently lower costs than the other. This happens tacitly, meaning that no information about the existence or about the strategy followed by the other agent is part of the information available either in the market or in each agent’s memory. Notice that predatory strategies are still consistent with the definition of Pareto efficiency but are not a proper collusion since they do differ from the Pareto-optimal $IS$, even if they appear to overlap in Figure 3.
Figure 4: Optimal execution strategies for 20 testing runs of 2500 iterations, using $\sigma=10^{-3}$.
Looking at the average strategies implemented by the agents in each testing run (Figure 4), it can be noticed how similarly as in the zero noise case, roughly speaking there is almost always one agent that trades faster than the other and again as in the previous case, the one agent that trades slower is the one with an higher $IS$. The comparison of Figure 4 and Figure 3shows that the agent that ‘predates’ the other is a fast trader and obtains lower costs of execution. This behaviour is more pronounced in this case, and we postulate that this is due to the increased level of noise for this experiment.
We conclude once again that, even if no explicit information about the trading strategies is shared by the agents, during the repeated game they are able to extrapolate information on the policy pursued by the competitor through the changes in the mid-price triggered by their own trading and by the other agent’s trading. Thus, it seems plausible that RL-agents modelled in this way are able to extrapolate information on the competitor’s policy and tacitly interact through either a collusive behaviour that leads to lower costs, either for both or for just one of them, resulting in a predatory behaviour or through a predatory behaviour where the costs of one agent are consistently higher than those of the other. Generally speaking, the set of solutions achieved is in line with the definition of the Pareto efficient set of solutions.
@@ 4.3 The large noise case
Finally, we study the interactions between the agents when the volatility of the asset is large. This corresponds to a relatively liquid market, since price changes are severely affected by the volatility level. Looking at Figure 5, we notice that in this market setup the higher level of volatility significantly affects the distribution of the $IS$ per iteration. In fact, the points in the scatter plot now distribute obliquely, even if for the most part they still lie in the second, third, and fourth quadrants. Because of the higher volatility level, very low $IS$ values for both agents might be attained per iteration and the structure of the costs per iteration is completely different with respect to to the previous cases. However, the distribution of the average $IS$s per testing run, i.e. the positions of the centroids, is still concentrated for the greatest part in the collusive strategies area in between the Pareto-optimum and the Nash equilibrium costs.
Figure 5: Scatter plot of the IS of the two agents for 20 testing runs of 2500 iterations in the large noise case ($\sigma=10^{-2}$).
Figure 6 shows the trading strategies for the agents. It can still be noticed how in general if one agent is more aggressive with its trading, the other tends to be not. This tacit behaviour of the agents resulting in faster and slower traders still takes place, and again the agent that is trading slower pays more in terms of $IS$, hence coupling the centroids’ distribution and the selling schedules in Figure 6 and Figure 5 reveals how a slower trading speed worsens the cost profile of one agent to the benefit of the other. The average trading strategy of the agents per testing run in the collusive cases basically revolves around a TWAP strategy that is in turn the Pareto-optimum for the considered problem.
Figure 6: Optimal execution strategies for 20 testing runs of 2500 iterations, using $\sigma=10^{-2}$.
@@ 4.4 Summary of results and comparison with the Pareto front
We have seen above that especially in the presence of significant volatility, the points corresponding to the iterations of a run tend to distribute quite widely in the scatter plot. The centroid summarises the average behaviour in a given run and provides a much more stable indication of the relation between the $IS$ of the two agents. To have a complete overview of the observed behaviour across the three volatility regimes, in Figure 7 we show in a scatter plot the position of the centroids of the $20\times 3$ testing runs. It can be seen that the centroids tend to concentrate in the collusive area, irrespective of the volatility level for the experiment. We notice how the agents’ costs tend to cluster in this area and to be close to the Pareto-optimal $IS$. To have a more detailed comparison of the simulation results with the theoretical formulation of the game, we numerically compute the Pareto-efficient set of solutions (or Pareto-front) [^6] The numerical Pareto front has been obtained using ‘Pymoo’ package in Python introduced in [pymoo]. of and represent it on the scatter plot. The Figure shows that, as expected, the centroids lie on the right of the front. More interestingly, they are mostly found between the front and the Nash equilibrium for all the levels of volatility and, in line with the definition of the Pareto-front, for an agent to get better the other has to be worse off. We further notice that the majority of points for the zero noise case lies very close the Pareto optimum or in the collusive area, thus the lower the volatility the easier it is to converge to the collusion strategy. In the other two cases we notice how, even roughly half of the centroids lie in the collusive area, the rest mostly lie not far from the numerical Pareto front.
Figure 7: Scatter plot of the IS centroids of the two agents in the $20\times 3$ testing runs of 2500 iterations each, for all the considered values of $\sigma$.
Figure 8: Average optimal execution strategy over the 20 testing runs of 2500 iterations, for all the values of $\sigma$ considered.
Figure 8 shows that, irrespective of the values considered for the volatility level $\sigma$, the average strategy does not differ much from the Pareto-optimum TWAP strategy. In the zero noise case, the average strategy of the two agents tends to be slightly more front loaded, but still lies between the Nash and the Pareto-optimum. As the $\sigma$ value increases, and thus in the moderate noise and large noise cases, we can see how both the agents tend to be less aggressive at the beginning of their execution in order to adopt a larger selling rate towards the end of their execution. This behaviour is stronger the higher the $\sigma$, due to the increased uncertainty brought by the asset volatility in addition to the price movements triggered by the trading of both the agents.
@@ 4.5 Variable volatility and misspecified dynamics
Our simulation results show that volatility plays an important role in determining the $IS$ in an iteration of the testing phase, although when averaging across the many iterations of a run, the results are more stable and consistent. In financial market, returns are known to be heteroscedastic, i.e. volatility varies with time. Thus, also from the practitioner’s point of view, it is interesting to study how agents, which are trained in an environment with a given level of volatility level, perform in a testing environment where the level of volatility is different. In particular, it is not a priori clear whether a collusive relationship between the two agents would still naturally arise when the price dynamics is different in the testing and in the training phase. Moreover, it is interesting to study how a change in the environment would impact the overall selling schedule and the corresponding costs. In the following, we study the agents’ behaviour in extreme cases, in order to better appreciate their behaviour under time-varying conditions. Specifically, using the same impact parameters as above in both phases, we learn the the weights of the DDQN algorithm in a training setting with $\sigma=10^{-9}$ and then we employ them in a testing environment where $\sigma=10^{-2}$. We then repeat the experiment in the opposite case with the volatility parameters switched, i.e. training with $\sigma=10^{-2}$ and testing with $\sigma=10^{-9}$.As before, we run $10$ testing runs of $2,500$ iterations each.
Figure 9: IS scatter for $20$ testing runs of $2500$ iterations, training $\sigma=10^{-9}$ and testing $\sigma=10^{-2}$.
Figure 10: Optimal execution strategies for $10$ testing runs of $2500$ iterations, training $\sigma=10^{-9}$ and testing $\sigma=10^{-2}$.
@@ 4.5.1 Training with zero noise and testing with large noise
When the agents are trained in an environment where $\sigma=10^{-9}$ and the weights of this training are used in a testing environment with $\sigma=10^{-2}$, we find that the distribution of the $IS$s per iteration is similar to the one that would be obtained in a both testing and training with $\sigma=10^{-2}$ scenario (see Figure 9). The centroids are still mostly distributed in the second, third and fourth quadrant, although some iterations and even a centroid, might end up in the first quadrant as in the $\sigma=10^{-9}$ case. In general, it can be seen how the centroids mostly lie in the collusive area of the graph, i.e in the square between the Pareto-optimal $IS$ and the Nash equilibrium, pointing out at the fact that collusive behaviours are still attainable even when the training dynamics are misspecified with respect to the testing ones.
@@ 4.5.2 Training with large noise and testing with zero noise
Figure 11: IS scatter for $20$ testing runs of $2500$ iterations, training $\sigma=10^{-2}$ and testing $\sigma=10^{-9}$.
@@
Figure 12: Optimal execution strategies for $10$ testing runs of $2500$ iterations, training $\sigma=10^{-2}$ and testing $\sigma=10^{-9}$.
@@
Figure 11 shows the result of the experiment where agents are trained in an environment where $\sigma=10^{-2}$ and tested in a market with $\sigma=10^{-9}$. We observe that the $IS$s are distributed in a way similar to the case where the agents were both tested and trained in an environment with $\sigma=10^{-9}$. Still, the centroids lie in all but the first quadrant, and the majority of them lies in the collusive area, showing how even with misspecified dynamics, a collusive behaviour naturally arises in this game. The trading schedules adopted show the same kind of fast-slow trading behaviour between the agents, where still the slower trader has to pay higher costs in terms of $IS$ (see Figure 12).
@@
Finally, irrespective of the encountered levels of volatility, it seems that what is more important is the market environment experienced during the training phase, thus the selling schedule implemented in a high volatility scenario with DDQN weights coming from a low volatility scenario are remarkably similar to the one found with both training and testing with $\sigma=10^{-9}$, and vice versa when $\sigma=10^{-2}$ is used in training and the low volatility is encountered in the test. It might be concluded that once that the agents learn how to adopt a collusive behaviour in one volatility regime, they are able to still adopt a collusive behaviour when dealing with another volatility regime.
@@ 5 Conclusions
In this paper we have studied how collusive strategies arise in an open loop two players’ optimal execution game. We first have introduced the concept of collusive Pareto optima, that is, a vector of selling strategies whose IS is not dominated by the IS of other strategies for each iteration of the game. We furthermore showed how a Pareto-efficient set of solutions for this game exists and can be obtained as a solution to a multi-objective minimisation problem. Finally, we have shown how the Pareto-optimal strategy, that is indeed the collusion strategy in this setup, is the TWAP for risk-neutral agents.
@@
As our main contribution, we have developed a Double Deep Q-Learning algorithm where two agents are trained using RL. The agents were trained and tested over several different scenarios where they learn how to optimally liquidate a position in the presence of the other agent and then, in the testing phase, they deploy their strategy leveraging on what learnt in the previous phase. The different scenarios, large, moderate and low volatility, helped us to shed light on how the trading interactions on the same asset made by two different agents, who are not aware of the other competitor, give rise to collusive strategies, i.e. strategies with a cost lower than a Nash equilibrium but higher than a proper collusion. This, in turn, is due to agents who keep trading at a different speed, thus adjusting the speed of their trading based on what the other agent is doing. Agents do not interact directly and are not aware of the other agent’s trading activity, thus strategies are learnt from the information coming from the impact on the asset price in a model-agnostic fashion.
@@
Finally, we have studied how the agents interact when the volatility parameter in the training part is misspecified with respect to the one observed by the agents in the testing part. It turns out that the existence of collusive strategies are still arising, and thus robust with respect to settings when models parameters are time varying.
@@
There are several possible extensions of our work. One obvious extension is to the setting where more than two agents are present and/or where more assets are liquidated, leading to a multi-asset and multi-agent market impact games, leveraging on the work done in [cordoni2024instabilities]. Second, impact parameters are constant, while the problem becomes more interesting when liquidity is time-varying (see [macrì2024reinforcement] for RL optimal execution with one agent). Third, we have considered a linear impact model, whereas it is known that impact coefficients are usually non-linear and might follow non-trivial intraday dynamics. Finally, the Almgren-Chriss model postulates a permanent and fixed impact, while many empirical results point toward its transient nature (see [bouchaud2009markets]). Interestingly, in this setting the Nash equilibrium of the market impact game shows price instabilities in some regime of parameters (see [schied2019market]), which are similar to market manipulations. It could be interesting to study if different market manipulation practices might arise when agents are trained, as in this paper, with RL techniques. The answer would be certainly of great interest also to regulators and supervising authorities.

### v2

@@ 4.2 The moderate noise case
In this setting, the mid-price dynamics is influenced both by the trading of the agents and by the volatility. The first thing that can be noticed in Figure 3 is that the $IS$ per iteration distributes almost only in the second, third and fourth quadrant of the plot. Moreover, the distribution of the points suggests again an inverse relation between the $IS$s of the two agents, i.e. a lower $IS$ of an agent is typically associated with a worsening of the $IS$ of the other agent. The centroids distribute accordingly. In fact, the centroids that lie in the supra-competitive area of the graph are as numerous as those that lie outside of that area on either the second or fourth quadrant. It can be seen that, outside the green rectangle, the agents tend to behave in a predatory way, meaning that one agent has consistently lower costs than the other. This happens tacitly, meaning that no information about the existence or about the strategy followed by the other agent is part of the information available either in the market or in each agent’s memory. Notice that predatory strategies are still consistent with the definition of Pareto efficiency but are not a proper collusion since they do differ from the Pareto-optimal $IS$, even if they appear to overlap in Figure 3.
Figure 4: Optimal execution strategies for 20 testing runs of $2,500$ iterations, using $\sigma=10^{-3}$.
Looking at the average strategies implemented by the agents in each testing run (Figure 4), it can be noticed how as in the zero noise case, roughly speaking there is almost always one agent that trades faster than the other and, again as in the previous case, the one agent that trades slower is the one with a higher $IS$. The comparison of Figure 4 and Figure 3 shows that the agent that ‘predates’ the other is a fast trader and obtains lower costs of execution. This behaviour is more pronounced in this case, and we postulate that this is due to the increased level of noise for this experiment.
We conclude once again that, even if no explicit information about the trading strategies is shared by the agents, during the repeated game they are able to extrapolate information on the policy pursued by the competitor through the changes in the mid-price triggered by their own trading and by the other agent’s trading. Thus, it seems plausible that RL-agents modelled in this way are able to extrapolate information on the competitor’s policy and tacitly interact through either a plausibly cooperative behaviour that leads to supra-competitive costs or through a predatory behaviour where the costs of one agent are consistently higher than those of the other. Generally speaking, the set of solutions achieved is in line with the definition of the Pareto efficient set of solutions.
@@ 4.3 The large noise case
Finally, we study the interactions between the agents when the volatility of the asset is large. This corresponds to a relatively liquid market, since price changes are severely affected by the volatility level and the price impact triggered by the agents is low in comparison. Looking at Figure 5, we notice that in this market setup the higher level of volatility significantly affects the distribution of the $IS$ per iteration. In fact, the points in the scatter plot now distribute obliquely, even if for the most part they still lie in the second, third, and fourth quadrants. Because of the higher volatility level, very low $IS$ values for both agents might be attained per iteration and the structure of the costs per iteration is completely different with respect to to the previous two cases. However, the distribution of the average $IS$s per testing run, i.e. the positions of the centroids, is still concentrated for the greatest part in the strategies’ area in between the Pareto-optimum and the Nash equilibrium costs.
Figure 5: Scatter plot of the IS of the two agents for 20 testing runs of $2,500$ iterations in the large noise case ($\sigma=10^{-2}$).
Figure 6 shows the trading strategies for the agents. It can still be noticed how in general if one agent is more aggressive with its trading, the other tends to be not. This behaviour of the agents, resulting in faster and slower traders, still takes place and again the agent that is trading slower pays more in terms of $IS$. Comparing the centroids’ distribution and the selling schedules in Figure 6 and Figure 5 reveals how a slower trading speed worsens the cost profile of one agent to the benefit of the other. The average trading strategy of the agents per testing run in the supra-competitive cases basically revolves around a TWAP strategy that is in turn the Pareto-optimum for the problem considered.
Figure 6: Optimal execution strategies for 20 testing runs of $2,500$ iterations, using $\sigma=10^{-2}$.
@@ 4.4 Summary of results and comparison with the Pareto front
We have seen above that especially in the presence of significant volatility, the points corresponding to the iterations in a run tend to distribute quite widely in the scatter plot. The centroid summarises the average behaviour in a given run and provides a much more stable indication of the relation between the $IS$ of the two agents. To have a complete overview of the observed behaviour across the three volatility regimes, in Figure 7 we show in a scatter plot the position of the centroids of the $20\times 3$ testing runs. It can be seen that the centroids tend to concentrate in the supra-competitive area, irrespective of the volatility level for the experiment. We notice how the agents’ costs tend to cluster in this area and to be close to the Pareto-optimal $IS$. To have a more detailed comparison of the simulation results with the theoretical formulation of the game, we numerically compute the Pareto-efficient set of solutions (or Pareto-front) [^10] The numerical Pareto front has been obtained using ‘Pymoo’ package in Python introduced in [undeff]. and represent it on the scatter plot. The Figure shows that, as expected, the centroids lie on the top-right of the front. More interestingly, they are mostly found between the front and the Nash equilibrium for all the levels of volatility and, in line with the definition of the Pareto-front, for an agent to get better the other has to be worse off. We further notice that the majority of points for the zero noise case lies very close the Pareto optimum or in the supra-competitive area, thus the lower the volatility the easier it is to converge to a supra-competitive strategy. In the other two cases we notice how, even roughly half of the centroids lie in the supra-competitive area, the rest mostly lie near the numerical Pareto front.
Figure 7: Scatter plot of the IS centroids of the two agents in the $20\times 3$ testing runs of $2,500$ iterations each, for all the considered values of $\sigma$.
Figure 8: Average optimal execution strategy over the 20 testing runs of $2,500$ iterations, for all the values of $\sigma$ considered.
Figure 8 shows that, irrespective of the values considered for the volatility level $\sigma$, the average strategy does not differ much from the Pareto-optimum TWAP strategy. In the zero noise case, the average strategy of the two agents tends to be slightly more front loaded, i.e. they tend to liquidate more at the beginning of the trading schedule, but still lies between the Nash and the Pareto-optimum. As the $\sigma$ value increases, and thus in the moderate noise and large noise cases, we can see how both the agents tend to be less aggressive at the beginning of their execution in order to adopt a larger selling rate towards the end of their execution. This behaviour is stronger the higher the $\sigma$, due to the increased uncertainty brought by the asset volatility in addition to the price movements triggered by the trading of both the agents.
@@ 4.5 Variable volatility and misspecified dynamics
Our simulation results show that volatility plays an important role in determining the $IS$ in an episode of the testing phase, although when averaging across the many iterations of a run, the results are more stable and consistent. In financial market, returns are known to be heteroscedastic, i.e. volatility varies with time. Thus, also from the practitioner’s point of view, it is interesting to study how agents, which are trained in an environment with a given level of volatility level, perform in a testing environment where the level of volatility is different. In particular, it is not a priori clear what behaviour among the two agents would naturally arise when the price dynamics is different in the testing and in the training phase. Moreover, it is interesting to study how a change in the environment would impact the overall selling schedule and the corresponding costs. In the following, we study the agents’ behaviour in extreme cases, in order to better appreciate their behaviour under time-varying conditions. Specifically, using the same impact parameters as above in both phases, we learn the the weights of the DDQN algorithm in a training setting with $\sigma=10^{-9}$ and then we employ them in a testing environment where $\sigma=10^{-2}$. We then repeat the experiment in the opposite case with the volatility parameters switched, i.e. training with $\sigma=10^{-2}$ and testing with $\sigma=10^{-9}$. We now run $10$ testing runs of $2,500$ iterations each.
Figure 9: IS scatter for 2020 testing runs of $2,500$ iterations, training $\sigma=10^{-9}$ and testing $\sigma=10^{-2}$.
Figure 10: Optimal execution strategies for $10$ testing runs of $2,500$ iterations, training $\sigma=10^{-9}$ and testing $\sigma=10^{-2}$.
@@ 4.5.1 Training with zero noise and testing with large noise
When the agents are trained in an environment where $\sigma=10^{-9}$ and the weights of this training are used in a testing environment with $\sigma=10^{-2}$, we find that the distribution of the $IS$s per iteration is similar to the one that would be obtained in a both testing and training scenario with $\sigma=10^{-2}$ (see Figure 9). The centroids are still mostly distributed in the second, third and fourth quadrant, although some iterations and even a centroid, might end up in the first quadrant as in the $\sigma=10^{-9}$ case. In general, it can be seen how the centroids mostly lie in the, supra-competitive area of the graph, i.e in the square between the Pareto-optimal $IS$ and the Nash equilibrium, pointing out at the fact that supra-competitive cost outcomes and strategies are still attainable even when the training dynamics are misspecified with respect to the testing ones.
@@ 4.5.2 Training with large noise and testing with zero noise
Figure 11: IS scatter for 2020 testing runs of $2,500$ iterations, training $\sigma=10^{-2}$ and testing $\sigma=10^{-9}$.
@@
Figure 12: Optimal execution strategies for $10$ testing runs of $2,500$ iterations, training $\sigma=10^{-2}$ and testing $\sigma=10^{-9}$.
@@
Figure 11 shows the result of the experiment where agents are trained in an environment where $\sigma=10^{-2}$ and tested in a market with $\sigma=10^{-9}$. We observe that the $IS$s are distributed in a way similar to the case where the agents were both tested and trained in an environment with $\sigma=10^{-9}$. Still, the centroids lie in all but the first quadrant, and the majority of them lies in the supra-competitive area, showing how even with misspecified dynamics, supra-competitive strategies naturally arise in this game. The trading schedules adopted show the same kind of fast-slow trading behaviour between the agents, where still the slower trader has to pay higher costs in terms of $IS$ (see Figure 12).
@@
Finally, irrespective of the encountered levels of volatility, it seems that what is more important is the market environment experienced during the training phase, thus the selling schedule implemented in a high volatility scenario with DDQN weights coming from a low volatility scenario are remarkably similar to the one found with both training and testing with $\sigma=10^{-9}$, and vice versa when $\sigma=10^{-2}$ is used in training and the low volatility is encountered in the test. It might be concluded that once that the agents learn how to adopt a supra-competitive behaviour in one volatility regime, they still behave supra-competitively even if they are dealing with another volatility regime.
5 Real tacit collusion or supra-competitive solution?
According to [undefab], <<Collusion is when firms use strategies that embody a reward-punishment scheme that rewards a firm for abiding by the supra-competitive outcome and punishes it for departing from it>>. Are we looking at real collusive behaviour then? Judging by the average trading strategies in Figure 8 one can infer that the agents have learnt to adopt a strategy compatible with a supra-competitive equilibrium, very close to the Pareto optimal strategy. Hence, from a ‘global’ point of view, the strategies suggest a cooperative behaviour by the agents that is directed at obtaining lower costs if compared to the Nash equilibrium strategy, as this strategy jointly minimises the costs for both the agents collectively. But this alone is not sufficient to talk about collusion, although tacit.
If we consider a more ‘local’ point of view, run-wise we see how the average of the testing iterations results in centroids that do distribute along the Pareto efficient set of solutions as in Figure 7, and this depends on the strategies that, as we saw, show different speeds of trading per agent. In fact, comparing the centroids in Figure 3 with the average trading patterns in Figure 4, it appears that the points in upper (lower) branch of the Pareto front correspond to cases where agent 1 trades slowly (fast) achieving a higher (lower) cost, while the opposite is true for agent 2. Thus, one can conjecture that the different speeds at which the agents trade in each iteration are a response to that deviation from the Pareto-optimal strategy, as they lie over the efficient frontier that is the set of solutions to the multi-objective optimisation problem.
Is it possible to interpret these deviations from the collusive strategy as a form of punishment? In our view, there is no simple answer to this question. On the one hand, considering the peculiar nature of the game itself, which is repeated and iterated because of the learning process, it is not possible to look at short deviations from the equilibrium and an eventual reversion to supra-competitive strategies, in the long run, because of a punishment mechanism. On the other hand, one can interpret the long run as a repetition of what was learned and observed in terms of the average strategy, which is in fact supra-competitive, as seen in Figure 8. In summary, we used the definition of collusion framed in [undefs]. However, in order to be conservative we cannot conclude that the observed behaviour is a tacit collusion in the sense of [undefab]. Clearly more analyses are required to answer this question, possibly considering longer iterations allowing to identify evidences of punishment by shocking the action of one of the agent as in [undefk]. We leave this interesting investigation for future research.
As a final note, we highlight that such behaviour could cause welfare reduction for market participants, in terms of a distorted price formation process.
In fact, it is interesting to notice that the observed trading behaviour might have adverse effects on other market participants’ welfare. In the Nash equilibrium the two agents trade more at the beginning of the time interval, the way information is incorporated into the price is relatively fast. On the contrary, when the two agents are colluding they spread more evenly the execution over the time window cooperatively. In this way, private information is incorporated more slowly. Ceteris paribus, this has a negative impact on other market participants by distorting the price formation process, since the mid-price would not accurately describe the market demand and offer mechanism for the stock under execution. This would negatively affect the price formation process for other traders and eventually reduce the welfare of traders who trade based on information available on the market.
In conclusion, even if there no direct evidence of a reward-punishment scheme adopted by the agents, there are possibly negative consequences in terms of welfare at the expense of other market participants, while the agents that are doing optimal execution would benefit by cost reductions.
@@ 6 Conclusions
In this paper we have studied how supra-competitive strategies arise in an open loop two players’ optimal execution game. We first have introduced the concept of collusive Pareto optima, that is, a vector of selling strategies whose $IS$ is not dominated by the $IS$s achievable by other strategies for each iteration of the game. We furthermore showed how a Pareto-efficient set of solutions for this game exists and can be obtained as a solution to a multi-objective minimisation problem. Finally, we have shown how the Pareto-optimal strategy, that is indeed the collusion strategy in this setup, is the TWAP for risk-neutral agents.
@@
As our main contribution, we have developed a Double Deep Q-Learning algorithm where two agents are trained using RL. The agents were trained and tested over several different scenarios where they learn how to optimally liquidate a position in the presence of the other agent and then, in the testing phase, they deploy their strategy leveraging on what learnt in the previous phase. The different scenarios, large, moderate and low volatility, helped us to shed light on how the trading interactions on the same asset made by two different agents, who are not aware of the other competitor, give rise to supra-competitive and Pareto-efficient strategies, i.e. strategies with a cost lower than a Nash equilibrium but higher than a collusion. This, in turn, is due to agents who keep trading at a different speed, thus adjusting the speed of their trading based on what the other agent is doing. Agents do not interact directly and are not aware of the other agent’s trading activity, thus strategies are learnt from the information coming from the impact on the asset price in a model-agnostic fashion.
@@
Finally, we have studied how the agents interact when the volatility parameter in the training part is misspecified with respect to the one observed by the agents in the testing part. It turns out that the existence of supra-competitive strategies possibly consistent with a tacit collusive behaviour is still present, and thus robust with respect to settings when models parameters are time varying.
@@
There are several possible extensions of our work. An obvious extension is to the setting where more than two agents are present and/or where more assets are liquidated, leading to a multi-asset and multi-agent market impact games, leveraging on the work done in [undeft]. Second, impact parameters are constant, while the problem becomes more interesting when liquidity is time-varying (see [undefaf] for RL optimal execution with one agent). Third, we have considered a linear impact model, whereas it is known that impact coefficients are usually non-linear and might follow non-trivial intraday dynamics. Finally, the Almgren-Chriss model postulates a permanent and fixed impact, while many empirical results point toward its transient nature (see [undefg]). Interestingly, in this setting the Nash equilibrium of the market impact game shows price instabilities in some regime of parameters (see [undefal]), which are similar to market manipulations. It could be interesting to study if different market manipulation practices might arise when agents are trained, as in this paper, with RL techniques. The answer would certainly also be of great interest to regulators and supervising authorities.